OpenStudy (anonymous):

You take out a 1-year, $5,000 loan at 8% simple annual interest. The lender charges you a $100 fee. Thinking of the fee as additional interest, what is the actual annual interest rate r you will pay?

OpenStudy (anonymous):

interest equals 5000 x 1 x 9/100 =$450 fee=250 total=700 rate=total*100 ___________=700*100/5000=70/5=14% 5000*1